1a. If the market's labor demand and supply functions were instead given by of LD=260-w and labor supply function of LS=40+10w, what would the equilibrium wage be? (Answer format is 35)

1b. What would be the equilibrium labor quantity?

1c. The government again imposes a payroll tax of 10%, and we will again think in terms of the wage paid by employers. What is the new equilibrium wage?

(Answer format is 35.8) Round to 1 decimal place, if necessary.

1d. What is the new equilibrium labor quantity used?
Answer #1

1.

a.

Equilibrium:

LD = LS

260 - w = 40 + 10w

220 = 11w

Equilibrium wage = 20.

b.

Equilibrium quantity = L = 40 + 10(20) = 240.

c.

With payroll tax of 10%:

Labor demand = LD = 260 - w(1.10)

Labor Supply: LS = 40 + 10w

Equilibrium after payroll tax:

LD = LS

260 - 1.1w = 40 + 10w

220 = 11.1w

Equilibrium wage = 19.82.

d.

Labor used = 40 + 10(19.82) = 238.2.

If you are satisfied with the answer, please provide a positive rating. Feel free to comment in case of queries.

Have a nice day ahead!